Session: Feb 2014 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Badary (Zadrutsky at Tue Mar 11 17:01:38 2014) Success:

Noto (m.paterno at Tue Mar 11 17:04:20 2014) Success:

Zelenchukskaya (Ptitcin A. at Tue Mar 11 17:04:57 2014) Success:

Svetloe (Zborovskiy A. at Tue Mar 11 17:05:13 2014) Success:

Medicina (S. Mariotti at Tue Mar 11 17:10:13 2014) Minor failures: RECO
MK5 restarted twice at the begin of the schedule

Yebes (Daniel Valero at Tue Mar 11 17:17:50 2014) Minor failures: LATE

Metsahovi (Minttu Uunila at Tue Mar 11 17:42:50 2014) Minor failures: TSYS
Negative Tsys values, ifa and ifb values normal, but strange values for individual bbc's

Urumqi (Peng Li at Tue Mar 11 23:53:59 2014) Minor failures: HIGH

Arecibo (KVN at Wed Mar 12 00:36:37 2014) Success:

Arecibo (KVN at Wed Mar 12 00:36:41 2014) Success:

Hartebeesthoek (Jonathan Quick at Wed Mar 12 07:58:47 2014) Minor failures: HIGH
No known problems apart from heavy rain.

Torun (K. Borkowski at Wed Mar 12 10:05:05 2014) Minor failures: ANT FS
Problems with antenna control occurred twice, at 15:50 and 16:26 (both solved within 10 minutes), which resulted in lack of fringes in the third ftp scan correlated. Also BBC07 was out of order apparently throughout entire experiment (moreover, weaker fringes for BBC01).

Onsala (Jun Yang at Wed Mar 12 12:11:25 2014) Success:
Thanks to the great weather (blue sky), Onsala sensitivity was quite good.

Effelsberg (Georgi at Wed Mar 19 11:52:50 2014) Success: